p 7-9 Allen Company and Barker Company are competitors in the same industry. Selecte financial data from their 2011 statements follow. Balance Sheet December 31, 2011 Barker Company 35,000 120,000 190,000 100,000 20,000 520,000 $985,000 Allen Company Cash Accounts receivable Inventory Investments Intangibles Property, plant, and equipment 10,000 45,000 70,000 40,000 11,000 180,000 $356,000 Total assets
ng-Term Debt-Paying Ability (P 7-9 coNTINUED) Allen Company Barker Accounts payable Bonds payable Preferred stock, $1 par Common stock, $10 par Retained earnings 60,000 100,000 50,000 100,000 46,000 $356,000 $165,000 410,000 30,000 280,000 100,000 $985,000 Total liabilities and capital ncome Statement For the Year Ended December 31, 2011 Allen Company Barker Company Sales Cost of goods sold Selling and administrative expenses Interest expense Income taxes Net income $1,050,000 725,000 230,000 10,000 42,000 $ 43,000 $2,800,000 2,050,000 580,000 32,000 65,000 S 73.000 Industry Averages: Times interest earned Debt ratic Debt/equity Debt to tangible net worth 7.2 times 40.3% 66.6% 72.7% Required Compute the following ratios for each company: 1. Times interest earned 2. Debt ratio 3. Debt/equity ratio 4. Debt to tangible net worth a. Which company has the better long-term debt position? Explain. to take on additional long-term debt? Explain c.
Answer #1

a . Following are the ratios :-

1. Time interest earned = Earning before interest and tax (EBIT) / interest expense

Statement showing calculation of EBIT of both the company is as follows :-

Allen company Barker company
Sales 1050000 2800000
(-) cost of goods sold (725000) (2050000)
(-) selling & administrative expense (230000) (580000)
Earning before interest & tax (EBIT) $95000 $170000

Time interest earned :

For Allen company = 95000/10000 = 9.5

For Barker company = 170000/32000 = 5.31

2 . Debt ratio = total liabilities / total assets

For Allen company = (60000+100000)/356000 = 0.45

For Barker company = (165000+410000)/985000 = 0.58

3 . Debt / equity ratio = total liability / shareholder's equity

For Allen company = 160000/196000 = 0.82

For Barker company = 575000/410000 = 1.40

4 . Debt Sent to tangible net worth = total liabilities / tangible net worth

Tangible net worth = total assets - total liabilities - intangible assets

Statement showing calculation of tangible net worth of both the company are :-

Particulars Allen company Barker company
Total assets 356000 985000
(-) total liabilities (160000) (575000)
(-) intangible asset (11000) (20000)
Tangible net worth $185000 $390000

For Allen company = 160000/185000 = 0.86

For Barker company = 575000/390000 = 1.47

b . Barker company's debt ratio is 0.58 which means approximately 58% of asset is financed by debt more than 50% is not that good sign. Further debt equity ratio of Barker company is 1.40 which goes to show that for every $1.40 of debt there is only $1 of equity to give support which is again not a good situation and it has an excessive debt position already . Also debt to tangible net worth of Barker company is 1.47 which is not good. Considering all these we can say already Barker company has excessive debt in it's balance sheet and it would not be advisable for it to take further debt. Therefore, Barker company is not in a position to take additional long term debt

c. Debt equity ratio is Allen company is 0.82 which means for every $1 equity there is only $0.82 debt which means creditors of the company are secured enough whereas debt equity ratio of Barker company is 1.40 which means for every $1 of equity there is $1.4 debt which means creditors of the company are not secured and not covered by the support of equity. Further debt ratio of Allen company is 0.45 whereas Barker company has 0.58 which means more assets of Barker company are financed by debt and they have more right on them . Further debt to Net tangible net worth of Barker company is 1.47 compared to that of alllen company which has 0.86 . Again Allen ompany has better ratios.

Considering all the ratios and their respective discussion we can say that allen company has better long term debt position than Barker company.